OpenStudy (anonymous):

Help! The unit of a circle. I missed this class lesson so I need to understand.-precalc find the terminal point P(x,y) on the unit circle given value t= 5(pie)/3

OpenStudy (anonymous):

\[-1/2,\sqrt{3}/2\]

OpenStudy (anonymous):

how did you get that though?

OpenStudy (anonymous):

made another typo\[1/2,- \sqrt{3}/2\]

OpenStudy (anonymous):

made another typo\[1/2,- \sqrt{3}/2\]

OpenStudy (anonymous):

counted one pi/3 + 2 pi /3 + ... + 5 pi /3 and then used the reference angle pi/3 and used the appropiate 4 quadrant angle

OpenStudy (anonymous):

did you get that?

OpenStudy (anonymous):

\[2\pi - 5\pi/3 =\pi/3\] but you are in the 4th quadrant so you must attach the appropiate signs. Is this making any sense to you?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Fabulous! Don't miss any more classes!